OpenStudy (anonymous):

2x+4y=8 3x+5y=1 can anyone help i have to substitution to solve the system of equations.but i dont get the right answer plz hlep

pooja195 (pooja195):

Have you tried the addition method?

pooja195 (pooja195):

\[\frac{ 2x+4y=8 }{ 3x+5y=1 }\]

OpenStudy (josedavid):

what grade are uin

OpenStudy (anonymous):

GOSH @josedavid dude you idiot

OpenStudy (anonymous):

10th but flocked algebra

OpenStudy (anonymous):

flunked ya mean @shelbylynn

OpenStudy (josedavid):

cool but u are taking 9 grade class just asking

OpenStudy (rayman5766):

lol you know whats funny there are 2 Shelby i seen so for you and shlbydancer

OpenStudy (anonymous):

@josedavid DUDE leave her alone stop stalling with her Gosh dude

pooja195 (pooja195):

Can you guys let me help if you guys arent?

pooja195 (pooja195):

@shelbylynn have you used the addition method?

OpenStudy (josedavid):

u think that i like her i dont like her i just want to help her

OpenStudy (rayman5766):

also pooja why did you leave me

OpenStudy (anonymous):

yes is it 5x+9y=9

OpenStudy (crissy15):

tell me if it helps

pooja195 (pooja195):

You have to cancel out the Ys or Xs and MUST have opposites

OpenStudy (anonymous):

Did the directions say you have to use substitution?

OpenStudy (anonymous):

@pooja195 when i said i need help u never replied me

OpenStudy (anonymous):

yes and is it 1x=9 then you do 9/1

pooja195 (pooja195):

\[\frac{3(2x+4y=8) }{ -2(3x+5y=1) }\]

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

@TheSmartOne

OpenStudy (anonymous):

@pooja195 u said when i tired tell u told u but u never helped me

pooja195 (pooja195):

@hiamme please stop going to every post typing that i did help you you just didnt want to put in work u were looking for answers u cant help if u dont work with me

OpenStudy (rayman5766):

ok pooja need your help

OpenStudy (rayman5766):

i will tag you also sorry to intude like this bye

OpenStudy (anonymous):

@pooja195 stop i tired i had work but ur the one who is not helping and telling people try and tell me i did tried kay

pooja195 (pooja195):

@shelbylynn please close this and open another post im so sorry :(

TheSmartOne (thesmartone):

Please stop spamming this post.

pooja195 (pooja195):

huh??? that made no sense if you have a problem feel free to report me

OpenStudy (anonymous):

you shouldnt report peeps jus cuz they dont answer a question ok???? youre all spamming in a jerkish way anyway

OpenStudy (anonymous):

ok